OWG Meeting 18-Jan-2012
- Review of minutes from 7-Dec-2011 meeting
- Software review 7,8-Jun - Mark, David, Elliott
- Counting house installation - Hovanes, Elliott
- furniture, phones, telecom, networking, racks, control room consoles, etc.
- Progress on 12GeV planning lines - David, Hovanes, Elliott
- ConDB, DAQ config, code mgmt, unblocking, event display, scripting, run info archive, storage mgmt
- Configuration database - David
- Elog and Accelerator division projects - Graham, Elliott
- DAQ group activities
- CODA3 alpha release - Graham, Dave A
- top-level ECS and gui - Vardan
- simulated raw event file - Dave A
- event disentangler - Carl
- automated readout list - Bryan
- EPICS - Hovanes, Nerses
- discriminator readout
- alarm system
- Solenoid Controls Redesign - Yi, Elliott
- Magnet cold by Nov-2012
- PLC programming
- quench detector
- coil taps, fast/slow DAQ integration, Labview, CRio vs PXI, etc.
- FMEA and redundancy
- Solenoid Controls Redesign - Yi, Elliott
- Beam tests - Alex
- Controls test stand - Hovanes, Elliott
- FDC cosmic ray test - Beni, Elliott
- Midas Elog - Elliott
- Maintenance/modification elog
- Manual database
- Procedures database
- Electronics, trigger and DAQ status reports - Chris, Alex, Dave, Fernando, Ben, Ed, William, Bryan
- PyIrmis - release due in Feb
- Mantis Task Tracker
1:30 PM Wed 18-Jan-2012 CC F326
Hovanes noted that everyone should contact Fernando if they need electronics boards or equipment, and contact Hovanes if they need VME CPU's. He asks that you contact them at least three months in advance.
1:30 PM Wed 1-Feb-2012 CC F326
Present: Elliott W, David L, Hovanes E, Beni Z, Somon T, Carl T, Mark I, Bryan M, Chris C, Dave A.
The June software review will mainly be focused on the offline, but may cover some online topics. Stay tuned...
Counting House Installation
Hovanes and others reported that the UPS will arrive soon and installation is being scheduled, depends on when we officially get beneficial occupancy of the counting house. We likely will need a support frame under the UPS. About 19 racks are in the control room ready to be installed in the rack room. Power feeds are being designed. Networking, wireless and phones work in the office area, printer soon. Interim networking will be installed in the hall. We need to decide where we need clean vs UPS power.
12 GeV Project Lines
Elliott, David and Hovanes are working on completing the "Planning" lines in the 12GeV project schedule. Many documents are out or under review, most complete work on the corresponding line. The project ETC will be reviewed soon so the lines may change.
David presented thoughts on the confDb first presented by Dmitry and Alex (see his slides for details) and notes that he is not in full agreement with everything they propose. They of course noted that this was a first attempt and are open to suggestions for improvement. Everyone agreed there should be minimal duplication of information between the configuration, conditions and calibration databases. Questions were raised concerning where base information is held for our board inventory: configDB, separate electronics inventory DB, IRMIS, whatever.
Dave also presented some prototype GUI's that access the configDB based on Python/GTK+/Glade. He reported that development was relatively easy once he got over the not-particularly-steep initial learning-curve barrier.
CODA Run Configs
Dave and Elliott proposed using SVN to archive CODA run configuration files. A CODA configuration is a set of files in a directory hierarchy so SVN is a natural choice. The SVN revision number tags the file set.
Elliott reported this is on hold until Accelerator + four halls meet next month.
Dave A reported that CODA3 code is in an integration/testing stage now, prior to the alpha release in the next couple of weeks. The EVIOv4 preliminary version is out in C and Java, C++ not done yet. Glenn is trying to put together a single lab-wide single board computer purchasing plan.
Hovanes reported that the prototype driver for JLab discriminator readout is done, report is in the DocDB, along with another doc on Struck scaler readout (we purchased one board, driver only working on VxWorks at the moment). He is continuing to look at the EPICS alarm system "BEAST" from SNS. We are waiting on the meeting with Accelerator to learn more about archivers (i.e. DataBrowser/MyA integration).
- Hovanes ordered an 8-axis motor controller for the soon-to-be-purchased Newport Goniometer. A visitor from Yerevan may work on this.
- Elliott reported that the target date for cooling down the solenoid is still Nov-2012. Controls redesign waiting on George's FMEA (determine level of redundancy needed) and whether the NI PCX or CRIO fast/slow daq will work. The quench detector is understood and under control.
Beni needs three more TDC's, will contact Fernando.
Chris reported they are purchasing boards now and gearing up for full-crate testing.